OpenStudy (starfireandrobin):

HELP!!! PLEASE!! I'm very stupid, unfortunately, and I do not know how to answer this. The width of a rectangle is 8 cm less than its length. The perimeter of the rectangle is 52 cm. What is the width of the rectangle? 7 cm 9 cm 15 cm 17 cm

OpenStudy (starfireandrobin):

@mathmale Please help :(

OpenStudy (mhchen):

Perimeter = width + width + length + length :) 52cm = x + x + (x+8) + (x+8) "Width is 8cm less than length" 52cm = 4x + 16 Combining Like-terms (simplifying) 36 = 4x Subtraction Property of Equality 9 = x Division Property of Equality Answer: 9cm.

OpenStudy (starfireandrobin):

Where did the x come in on length times length??

OpenStudy (mathmale):

This being algebra, you need to represent unknown quantities with letters, such as x, y, z, P (for perimeter), A (for area), and so on.

OpenStudy (mathmale):

I'd recommend that you use letters such as l and w to represent length and width.

OpenStudy (starfireandrobin):

I know that, but I do not understand the (x + 8) + (x +8).

OpenStudy (starfireandrobin):

I would have thought it would solely be 8 + 8.

OpenStudy (mhchen):

Since the width is 8cm less than the length, I represented the width as x, and since the length is 8cm bigger than the width, I represented it as x+8.

OpenStudy (mathmale):

"The width of a rectangle is 8 cm less than its length:" If w represents the width and l the length, then w=l-8 [all in cm]. (Generally the width is less than the length.)

OpenStudy (starfireandrobin):

Can someone greatly explain this to me???

OpenStudy (starfireandrobin):

@mhchen @mathmale @Vuriffy

OpenStudy (starfireandrobin):

Even if I would have remembered the equation, P = 2w + 2l, I would not have thought of combining 4 X's and doing what @mhchen did

OpenStudy (starfireandrobin):

Please guys @mathmale @mhchen @Vuriffy

OpenStudy (starfireandrobin):

PLEASE I need to finish this I'm so tired

OpenStudy (starfireandrobin):

I'm so stupid that I can't even figure this out on my own

OpenStudy (mhchen):

Okay look, so the width is 8cm less than the length right? If we represented the width with x, then the length is x+8. Get it? OR by the same logic, we can make the width x-8 and the length as x. Same thing, got it? We just apply this to the formula you just wrote. P=2l+2w

OpenStudy (vuriffy):

Another way at going at this is by using that x+8 but plugging in values.

OpenStudy (vuriffy):

Let's try 7. (7+8)+(7+8)+(15-8)+(15-8) 44 = 52. Incorrect.

OpenStudy (starfireandrobin):

But if it's 9. then isn't it 9 + 9 = 18 and 8 + 8 = 16?? That's not 52

OpenStudy (starfireandrobin):

Or am I going about this totally wrong?

OpenStudy (vuriffy):

Let's try 9. (9+8)+(9+8)+(17-8)+(17-8) 17 + 17 + 9 + 9 52 = 52.

OpenStudy (vuriffy):

You may figure out an equation then plug in values from the answers to swiftly find what you are looking for.

OpenStudy (vuriffy):

A rectangle has 4 sides; two sides have the same value.

OpenStudy (starfireandrobin):

I don't get why we have to do 9 + 8 *2 and then 17 + 8 *2 I don't get it

OpenStudy (vuriffy):

As aforementioned, a rectangle has 4 sides. Two sides of the rectangle have the same value, and the other two sides of the rectangle have the same value.

OpenStudy (mathmale):

Earlier I suggested that you represent the width by l-8, since the width is "8 less than the length." The perimeter is given; it is 52 cm. The perimeter is the sum of 2 l and 2 w. Write this as an equation, please.

OpenStudy (vuriffy):

Which is why the value is there twice. (17 + 17) + (9 + 9)

OpenStudy (mathmale):

Once again: The formula for perimeter is P = 2 l + 2 w (where 'l" signifies length)